Jay, a rival of Dons, is the first to complete the 15 km time trial run of the competition. Jay is 75 kg and averages 4.4 m/s during the first half of the race. He then runs at a speed of 4.2 m/s until the last 200 meters, which he covers at 4.5 m/s. At what average speed must Don run to be at Jay?
Answer #1

Total time taken by jay to complete the race =

7500/4.4 + 7300/4.2 +200/4.5 = 3487.085 sec

Average speed = 15000/ 3487.085 = 4.3015 m/s

So don must run with this average speed or greater than this speed to win the race

Please rate it up thanks :)